OpenStudy (lainee03):

how do i find the absolute extremum of t sqr(4-t^2) [-1,2]

zepdrix (zepdrix):

Hi there! \(\Large \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (lainee03):

\[t \sqrt{4-t^2} \left[ -1,2 \right]\]

OpenStudy (lainee03):

Hi

zepdrix (zepdrix):

I guess to start, we would want to make sure that our function is `continuous` over the given interval. So in this case, our function has a domain of:\[\Large t\in [-2,2]\]So we don't have any problems there, good good.

OpenStudy (lainee03):

dont i have to sqr both sides and then take the derivative if each side

zepdrix (zepdrix):

\[\Large f(t)\quad=\quad t\sqrt{4-t^2}\]To find extrema, we'll take the derivative and set it equal to zero.\[\Large f'(t)\quad=\quad ?\] Mmm no let's not square each side! That will cause a big mess for us. Looks like we'll want to apply the product rule for this one.

zepdrix (zepdrix):

\[\Large f'(t)\quad=\quad \color{royalblue}{(t)'}\sqrt{4-t^2}+t\color{royalblue}{\left(\sqrt{4-t^2}\right)'}\]There is our setup for product rule. We need to take the derivative of the blue pieces. Make sense so far? :o

OpenStudy (lainee03):

right so it would be \[1\sqrt{4-t^2}+t(2t)\]

OpenStudy (lainee03):

is that right

zepdrix (zepdrix):

Hmm 2t?\[\Large \frac{d}{dx}\sqrt{4-t^2}\quad=\quad \frac{d}{dx}(4-t^2)^{1/2}\]Applying the power rule:\[\Large \frac{1}{2}(4-t^2)^{-1/2}\color{orangered}{\frac{d}{dx}(4-t^2)}\]We have to apply the chain rule, that's where the orange part is coming from:\[\Large =\frac{1}{2}(4-t^2)^{-1/2}\color{orangered}{(-2t)}\]If we simplify it down a bit, and rewrite our 1/2 power as a root,\[\Large =\frac{-t}{\sqrt{4-t^2}}\]

OpenStudy (lainee03):

ok so now all i do is substitute -1 and 2 in for t right

zepdrix (zepdrix):

So we found our derivative: \[\Large f'(t)\quad=\quad \sqrt{4-t^2}+\frac{-t^2}{\sqrt{4-t^2}}\]Now we want to set the derivative equal to zero and solve for t to find critical points.\[\Large 0\quad=\quad \sqrt{4-t^2}+\frac{-t^2}{\sqrt{4-t^2}}\]

zepdrix (zepdrix):

Start by getting a common denominator:\[\Large 0\quad=\quad \frac{(4-t^2)-t^2}{\sqrt{4-t^2}}\]

OpenStudy (lainee03):

how did you get t^2 on top

OpenStudy (lainee03):

i know that the common denominator is 4-t^2

zepdrix (zepdrix):

\[\Large f'(t)\quad=\quad \color{royalblue}{(t)'}\sqrt{4-t^2}+t\color{royalblue}{\left(\sqrt{4-t^2}\right)'}\]So our derivative gave us this:\[\Large f'(t)\quad=\quad \color{orangered}{(1)}\sqrt{4-t^2}+t\color{orangered}{\left(\frac{-t}{\sqrt{4-t^2}}\right)'}\]In the second term there, I brought the black t into the numerator and multiplied it by the -t.

OpenStudy (lainee03):

ok i see now .... i missed that "t"

zepdrix (zepdrix):

Woops I forgot to take the prime off of the orange block :) it shouldn't be there anymore.

OpenStudy (lainee03):

right i figured that...

zepdrix (zepdrix):

Normally, to solve for t, we would set the numerator equal to zero. To find critical points we will also have to consider places where the derivative might be undefined. For our problem that means we need to also set the denominator equal to zero.\[\Large 0\quad=\quad \sqrt{4-t^2}, \qquad\qquad\qquad 0\quad=\quad (4-t^2)-t^2\]So we need to solve for t each in each of these cases.

zepdrix (zepdrix):

What values do you get for t? :)

OpenStudy (lainee03):

-1,2

zepdrix (zepdrix):

Hmmm...

zepdrix (zepdrix):

For this one:\[\Large 0\quad=\quad \sqrt{4-t^2}\]Squaring each side:\[\Large 0\quad=\quad 4-t^2\]Solving for t gives us:\[\Large t\quad=\quad \pm \sqrt4\qquad\to\qquad t\quad=\quad\pm2\]

OpenStudy (lainee03):

right that is what i got for that one

zepdrix (zepdrix):

For the numerator we should get something similar: 2 answers due to the square.

zepdrix (zepdrix):

\[\Large 0\quad=\quad(4-t^2)-t^2\]\[\Large 0\quad=\quad 4-2t^2\]

zepdrix (zepdrix):

Hmm I don't think it'll give us -1 :O try again maybe? :D

OpenStudy (lainee03):

ok \[4-2t^2= 0 \rightarrow 2t^2=-4 \rightarrow 2t^2\div2 =4\div2\]

OpenStudy (lainee03):

that is how i am trying to solve this...am i doing it wrong?

zepdrix (zepdrix):

\[\Large t^2\quad=\quad 2\]Nope, looks good so far!

OpenStudy (lainee03):

ok so then t^2=2 would be 1 right

OpenStudy (lainee03):

if you squared it

zepdrix (zepdrix):

taking the square root of each side gives us:\[\Large t\quad=\quad \pm\sqrt2\]

OpenStudy (lainee03):

ok so our critical numbers are \[\pm 2,\pm \sqrt{2}\]

zepdrix (zepdrix):

Yes,\[\Large t=-2,\;-\sqrt2,\;\sqrt2,\;2\]Recall that our interval is \(\Large t\in[-1,2]\) Do all 4 of our critical points fall inside of this interval?

OpenStudy (lainee03):

no

zepdrix (zepdrix):

Hmm so we'll have to exclude any values that don't fall in our interval. Which I guess ends up leaving us with:\[\Large t=\sqrt2,\;2\]

OpenStudy (lainee03):

right

OpenStudy (lainee03):

so now we take our critical numbers and plug in

zepdrix (zepdrix):

So now comes the easier part. We have our `critical points` and our `end points`. We plug all of those values into the function f(t) and compare. The largest output is our max, the smallest our min.

zepdrix (zepdrix):

\[\Large f(-1)\quad=\quad ?\]\[\Large f(\sqrt2)\quad=\quad?\]\[\Large f(2)\quad=\quad ?\]

OpenStudy (lainee03):

so we plug into the original problem not the derivative...right

OpenStudy (lainee03):

ok so i got f(-1)=-2.24 f(\[\sqrt{2} =2\] f(2) = 0

zepdrix (zepdrix):

Mmm ok good! So x=-1 gave us the smallest value and x=sqrt2 gave us the largest value. So we've found our max/min points within the given interval!

zepdrix (zepdrix):

Write max/min as ordered pairs.

OpenStudy (lainee03):

max (\[\sqrt{2}\],2)

OpenStudy (lainee03):

min (-1,-2.24)

zepdrix (zepdrix):

yay good job \c:/